OpenStudy (anonymous):

Solve this system of equations using the addition method. x + y = 8 -x + 2y = 7

OpenStudy (ivettef365):

just add, and x will cancel right away

OpenStudy (ivettef365):

x + y = 8 -x +2y = 15 -------------

OpenStudy (anonymous):